Present Value for Various Compounding Periods Find the present value of $400 due in the future under each of the following conditions. Do not round intermediate calculations. Round your answers to the nearest cent.

4% nominal rate, semiannual compounding, discounted back 5 years

$ _______

4% nominal rate, quarterly compounding, discounted back 5 years

$________

4% nominal rate, monthly compounding, discounted back 1 year

$________
Answer #1

Answer:

(A) 4% nominal rate, semiannual compounding, discounted back 5 years

FV = $400

PV = (rate, nper, pmt, fv, type) = PV(4%/2, 10, 0, -400, 0) = $328.14

4% nominal rate, semiannual compounding, discounted back 5 years = $328.14

(B) 4% nominal rate, quarterly compounding, discounted back 5 years

PV = (rate, nper, pmt, fv, type) = PV(4%/4, 20, 0, -400, 0) = $327.82

4% nominal rate, quarterly compounding, discounted back 5 years = $327.82

(C)  4% nominal rate, monthly compounding, discounted back 1 year

PV = (rate, nper, pmt, fv, type) = PV(4%/12, 12, 0, -400, 0) = $384.34

4% nominal rate, monthly compounding, discounted back 1 year is = $384.34